2. Country A’s current GDP is $1,000,000. It is growing at the rate of 2% per year. It has a current population of 2,500 which is growing at 3% a year.

(a) Using the rule of 70, how long will it be before Country A’s GDP doubles (round off to the nearest value)? What will it’s per-capita GDP be in that year?

(b) Using the rule of 70, how long will it be before Country A’s population doubles (round off to the nearest value)? What will it’s per-capita GDP be in that year?

3. Country B’s current GDP is $500,000. It is growing at the rate of 8% per year. It has a current population of 5,000 which is growing at 1.5% a year.

(a) Using the rule of 70, how long will it be before Country B’s GDP doubles (round off to the nearest value)? What will it’s per-capita GDP be in that year?

(b) Using the rule of 70, how long will it be before Country B’s population doubles (round off to the nearest value)? What will it’s per-capita GDP be in that year?

4. Approximately when will Country B’s per-capita income equal Country A’s? why?
Answer #1

Using Rule of 70,

Doubling period = 70 / Annual growth rate

(2)

(a)

GDP Doubling period = 70 / 2 = 35 years

Population after 35 years = 2,500 x (1.03)35 = 2,500 x 2.8139 = 7,034.66

GDP per capita after 35 years = GDP / Population = ($1,000,000 x 2) / 7,034.66 = $284.31

(b)

Population doubling period = 70 / 3 = 23 years

GDP after 23 years = $1,000,000 x (1.02)23 = $1,000,000 x 1.5769 = $1,576,900

GDP per capita after 23 years = GDP / Population = $1,576,900 / (2,500 x 2) = $315.38

NOTE: As per Answering Policy, 1st question is answered.